consider market in which demand and supply equations are given by the following equations.

Qd = 400- 20P

Qs = 40P - 200

calculate equilibrium price and quantity


1
Expert's answer
2021-11-21T17:10:54-0500

Qd = Qs

400- 20P = 40P - 200

600 = 60P

P= 600/60 = 10

Qd = 400 - 20P = 400 - 20*10 =400-200 = 200

Qs = 40P - 200 = 40*10 - 200 = 400-200 = 200


Conclusion:

Equilibrium P = 10

Equilibrium Q = 200